OpenStudy (anonymous):

Prove: Sin Theta-sin theta(cos^2 theta)=sin^3 theta. I have to show all work. I already know I have to factor out the sin theta from the left side and that I have to use the identity : sin^2 θ + cos^2 θ = 1but I'm not sure where to go from there

OpenStudy (jdoe0001):

\(\bf sin(\theta)-sin(\theta)cos^2(\theta)=sin^3(\theta)?\)

OpenStudy (jdoe0001):

hmmm

OpenStudy (anonymous):

yes. When I factor out sin theta i get sin theta(1-cos^2 theta)=sin^3 theta I have no idea where to go from there

OpenStudy (anonymous):

wait... how did you get that?

OpenStudy (jdoe0001):

hmmm ok so \(\bf {\color{brown}{ sin^2(\theta)+cos^2(\theta)=1\to cos^2(\theta)=1-sin^2(\theta)}} \\ \quad \\ \quad \\ sin(\theta)-sin(\theta)cos^2(\theta)=sin^3(\theta) \\ \quad \\ sin(\theta)-sin(\theta)[{\color{brown}{ 1-sin^2(\theta)}}]\implies ?\)

OpenStudy (anonymous):

do i distribute \[-\sin \theta [1-\sin^2 \theta]\]

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

\[\sin \theta-\sin \theta +\sin^3 \theta\]

OpenStudy (anonymous):

Can you subtract \[\sin^3 \theta \] from both sides? and make both sides =0?